# Mind the Gap: a Spectral Analysis of Rank Collapse and Signal Propagation in Transformers

### v1

@@ 2.1 Revisiting rank collapse & exploding gradients
Rank collapse.
As Theorem 2 reveals, for increasing $T$, a Random Markov matrix becomes effectively rank-one. We accurately describe the rate of this decay through the notion of stable rank, defined for any non-zero $\mathbf{M}\in\mathbb{R}^{m\times n}$ as
Naturally, any definition of “rank collapse” relies on a proxy for discrete rank, as the random matrices in question are almost surely full-rank, making it uninformative to refer to their actual rank. For example, [DCL21] consider the “one-infinity norm” of the residual (the difference between a matrix and its best approximation of the form $\mathbf{1}\mathbf{x}^{\top}$), defined as $\sqrt{\lVert\mathrm{res}(\mathbf{M})\rVert_{1}\lVert\mathrm{res}(\mathbf{M})\rVert_{\infty}}$, while [NAB+22] use $\sum_{i,j}(\mathbf{M}\mathbf{M}^{\top})_{i,j}$, which is maximised when all rows of $\mathbf{M}$ are identical. We choose stable rank as our preferred proxy due to its clear geometrical interpretation and simple definition in terms of singular values.
Given an isometric input $\mathbf{X_{0}}$ with $\boldsymbol{\Sigma}_{0}\coloneqq\mathbf{X}_{0}\mathbf{X}^{\top}_{0}=\mathbf{I}$, we are interested in understanding how the stable rank of the covariance matrix at layer $\ell$,
@@
Proposition 3 (Rank collapse in width).
@@
with overwhelming probability [^4] An event $E_{n}$ holds with overwhelming probability if, for every $A>0$, $\mathbb{P}(E_{n})\geq 1-C_{A}n^{-A}$, for some constant $C_{A}$. As the name suggests, $E_{n}$ is more likely to hold if it occurs with overwhelming probability than with high probability, as defined in [Tao12].. Moreover, the convergence happens at a polynomial rate, i.e.
@@
A well-known issue that can disrupt training across various neural network architectures is the vanishing or exploding of gradients; see [Han18]. For attention-only transformers with degenerate attention, [NAB+22] demonstrate that the gradients with respect to $\mathbf{W}^{V}_{\ell}$ vanish. Our model ((3)) allows for more general random attention while using a different scaling that makes the same quantity explode rather than vanish. Proposition  4 provides a lower bound on the rate at which the gradient grows.
@@
2.2 Attention without the gap
As previously seen, a Random Markov matrix $\mathbf{A}\in\mathbb{R}^{T\times T}$ can be written as
where $\mathbf{1}_{T\times T}\coloneqq(1,\cdots,1)^{\top}(1,\cdots,1)$ is the all-ones matrix and $\mathbf{A}^{\perp}$ has a limiting spectrum resembling that of a Gaussian matrix. Therefore, $\mathbf{A}$ is a rank-one perturbation of $\mathbf{A}^{\perp}$, whose spectral radius is $\lambda_{1}(\mathbf{A}^{\perp})=O(T^{-1/2})$. Although the rank-one perturbation $\frac{1}{T}\mathbf{1}_{T\times T}$ cannot disturb the bulk of the spectrum, it causes the largest eigenvalue to “escape” from the bulk to $1$, creating a spectral gap.
In light of this, we can slightly modify the attention mechanism to eliminate the outlier—and thus the gap—simply by replacing $\mathbf{A}$ with $\mathbf{A}^{\perp}$ at every layer, i.e.
@@ 2.2 Attention without the gap
where $\mathbf{A}^{\perp}_{\ell}\coloneqq\mathbf{A}_{\ell}-T^{-1}\mathbf{1}_{T\times T}$. Note that this modification is applied only to the attention matrices (and not to the signal representation) and $\mathbf{X}^{\perp}_{\ell}$ serves as shorthand for the signal at layer $\ell\geq 1$ in a network whose $\mathbf{A}_{\ell}$’s are replaced with $\mathbf{A}^{\perp}_{\ell}$’s as in (9). We set $\mathbf{X}^{\perp}_{0}=\mathbf{X}_{0}$.
Since the modified attention exhibits no spectral gap (see Lemma 3 in section A.2), the stable rank of the covariance matrix $\boldsymbol{\Sigma}^{\perp}_{\ell}\coloneqq\mathbf{X}^{\perp}_{\ell}{\mathbf{X}^{\perp}_{\ell}}^{\top}$ no longer collapses to 1 in width, as detailed in Proposition 5 (cf. Proposition 3).
@@
Let $\mathbf{X}^{\perp}_{\ell}=\mathbf{A}^{\perp}_{\ell}\mathbf{X}^{\perp}_{\ell-1}\mathbf{W}^{V}_{\ell}$ be the signal at layer $\ell$ in our modified model ((9)) and $\mathbf{\Sigma}^{\perp}_{\ell}\coloneqq\mathbf{X}^{\perp}_{\ell}{\mathbf{X}^{\perp}_{\ell}}^{\top}\in\mathbb{R}^{T\times T}$ be its covariance matrix. Then, almost surely, the rank does not collapse, i.e., there exists a constant $C_{\ell}>0$ such that,
@@
Let $\mathbf{X}^{\perp}_{\ell}=\mathbf{A}^{\perp}_{\ell}\mathbf{X}^{\perp}_{\ell-1}\mathbf{W}^{V}_{\ell}$ be the signal at layer $\ell$ in our modified model ((9)). Then, in expectation, the squared norm of the gradients grow linearly with $d$, i.e. there exists a constant $C>0$ such that,
@@
2.3 Can transformers achieve dynamical isometry?
So far, we have established (i) the existence of an outlier eigenvalue/singular value in the spectrum of softmax-based attention matrices, and (ii) that removing this outlier helps with rank collapse and exploding gradients. In the absence of the outlier, we can take a further step to analyse the bulk of the spectra of the network’s token-wise covariance and input-output Jacobian.
Let $\mathbf{X}^{\perp}_{\ell}=\mathbf{A}^{\perp}_{\ell}\mathbf{X}^{\perp}_{\ell-1}\mathbf{W}^{V}_{\ell}$ be the signal at layer $\ell$ in our modified model ((9)) and $\mathbf{\Sigma}^{\perp}_{\ell}=\mathbf{X}^{\perp}_{\ell}{\mathbf{X}^{\perp}_{\ell}}^{\top}\in\mathbb{R}^{T\times T}$ be its covariance matrix. Let the underlying Random Markov matrices $\mathbf{A}_{\ell}$ have variance $\sigma_{A}^{2}$ and $\mathbf{W}^{V}_{\ell}$ have i.i.d. $\mathcal{N}(0,\sigma_{V}^{2})$ entries. Let $\boldsymbol{\Sigma}^{\perp}_{0}=\mathbf{I}$ and $\mathcal{D}_{\ell}$ be the limiting singular value distribution of $\boldsymbol{\Sigma}^{\perp}_{\ell}$. Then the mean and variance of $Z\sim\mathcal{D}_{\ell}$ are given by
$\mathbb{E}(Z)$ $=(\sigma_{A}\sigma_{V}/\sqrt{\gamma})^{2\ell},$ (12)
where $\gamma\coloneqq\frac{T}{d}\in(0,1]$.
It is evident from the above proposition that simultaneously controlling both the mean and variance of $\mathcal{D}_{\ell}$ is not feasible. Model ((9)) does not have enough hyperparameters to achieve this balance. Indeed, to prevent the mean from growing or shrinking exponentially with depth, the product $\sigma_{A}\sigma_{V}$ must equal $\sqrt{\gamma}$. However, this constraint leads to the variance increasing linearly with $\ell$.
The goal is to ensure that the spectral energy of the Jacobian concentrates around 1, thereby minimising distortion of the input space geometry—a property often referred to as the dynamical isometry in the literature (see [PSG17]). For our model ((9)), it is straightforward to show
where $\otimes$ denotes the Kronecker product. Proposition  8 describes the moments of the limiting squared singular value distribution of $\mathbf{J}_{\ell}$.
Let $\mathbf{X}^{\perp}_{\ell}=\mathbf{A}^{\perp}_{\ell}\mathbf{X}^{\perp}_{\ell-1}\mathbf{W}^{V}_{\ell}$ be the signal at layer $\ell$ in our modified model ((9)). Let the underlying Random Markov matrices $\mathbf{A}_{\ell}$ have variance $\sigma_{A}^{2}$ and $\mathbf{W}^{V}_{\ell}$ have i.i.d. $\mathcal{N}(0,\sigma_{V}^{2})$ entries. Let $\mathcal{D}_{\ell}$ be the limiting distribution of the squared singular values of $\mathbf{J}_{\ell}\coloneqq\partial\mathbf{X}^{\perp}_{\ell}/\partial\mathbf{X}_{0}$ [^5] In a minor abuse of notation, we may write $\partial\mathbf{X}^{\perp}_{\ell}/\partial\mathbf{X}_{0}$ as a shorthand for $\partial\,\mathrm{vec}(\mathbf{X^{\perp}_{\ell}})/\partial\,\mathrm{vec}(\mathbf{X}_{0})$.. Then the mean and variance of $Z\sim\mathcal{D}_{\ell}$ are given by
$\mathbb{E}(Z)$ $=(\sigma_{A}\sigma_{V})^{2\ell},$ (17)
$\mathrm{Var}(Z)$ $=\ell(\ell+2)(\sigma_{A}\sigma_{V})^{4\ell}.$ (18)
3 Experiments
Rank collapse.
@@ 3 Experiments
As an input signal propagates through a transformer, we can address both forms of rank collapse—across width and depth—by eliminating the spectral gap induced by the attention matrix at each layer. Figures 2 and 4 reinforce our findings, showing that our removing the gap consistently mitigates rank collapse even in multi-layer transformers that include additional components such as LayerNorm, skip connections, or both. It is crucial to understand that rank collapse in depth is an inherent consequence of successive matrix multiplications. Therefore, architectural modifications can only slow the collapse rather than completely prevent it. We demonstrate this by showing that rank collapse in depth persists even when the attention matrix is set to the identity matrix—an extreme case with the highest possible stable rank and no spectral gap. Another possible way to slow down rank collapse in depth (though not eliminate it) is to set the value matrices as orthogonal matrices.
(a) Spectral gap implies rank collapse in width.
@@
Figure 2: Rank collapse occurs both in width and depth. At layer one, our fix effectively prevents the rank from collapsing in width. Although rank collapse in depth occurs regardless of the presence of the spectral gap, our fix consistently slows the collapse—a feat no other module achieves.
Exploding gradients.
After passing an isometric input through the network, we compute the gradient norm as defined in (6). While our theory establishes a lower bound on the gradient norm at layer $1$ that scales linearly with width, Figure 4 confirms an overall linear growth, supporting the order-optimality of our result. This linear trend persists even in more general settings that incorporate additional modules, such as LayerNorm or skip connections. Notably, the removal of the spectral gap affects only the slope of the gradient norm’s growth, mirroring the behaviour observed in the lower bound derived in equations (4) and (6) for the single-layer case.
When the attention is a Random Markov matrix, gradient norms are effectively controlled with depth by either applying LayerNorm or removing the spectral gap, as illustrated in Figure 5, where the derived lower bound of $T^{L-1}$ is confirmed. Shifting to the more complex case of key-query attention in a multi-layer network, Figure 5 also demonstrates that our theoretical lower bound, derived for Random Markov matrices, remains valid. Interestingly, this bound becomes looser as depth increases, indicating that gradient norm explosion occurs at an even faster rate than predicted by our analysis; see the complementary Figure 9 in the Appendix. Moreover, in the multi-layer case, removing the largest singular value alone is no longer sufficient to prevent exploding gradients, suggesting more complex dynamics around the singular values of the attention matrix. One potential explanation is that the key-query attention spectrum now includes multiple outliers, whereas our method only addresses a single one. This hypothesis is explored further in the remainder of this section; see Figure 7.
Figure 3: The rank collapses in width in the
first layer across architectures.
Figure 4: At layer one, the gradient norm
scales linearly with width.
Figure 5: In multi-layer transformers with Random Markov attention (left), our “remove the gap” fix is effective, as we can precisely address the single outlier in the spectrum. However, with conventional key-query attention (right), the spectra of the attention matrices become more complex with depth, often exhibiting multiple outlier eigenvalues. This increased complexity reduces the effectiveness of our fix in controlling the gradient norm, as it only targets a single outlier, leaving other gaps untouched.
Training.
We evaluate our “remove the gap” proposal on a task designed to learn the entrywise Heaviside function; see section A.3 for implementation details. While our theoretical analysis does not address training dynamics, the experiments still offer valuable insights. In Figure 6, we present examples of training several architectural variants with and without our “remove the gap” solution. Further large-scale experiments are necessary to assess potential training benefits since the provided simulations are inconclusive. Beyond the first layer, key-query attention matrices are not Random Markov matrices, therefore their spectral properties are not well-known. For instance, we observe the emergence of additional outliers in the spectrum across layers, as shown in Figure 7. Investigating the configuration of the bulk and outliers in this context could lead to a natural solution for signal propagation issues by eliminating all outliers—an insight derived from our analysis.
Figure 6: Examples of training loss curves with and without removing the gap for single-layer (top row) and two-layer (bottom row) transformers.
Figure 7: Entries (top row) of a $T\times T$ key-query attention matrix (with $T=100$), along with its spectrum (bottom row) across layers. We indicate the number of eigenvalues whose magnitudes exceed a threshold of $0.5$ to signify the presence of multiple outliers. The layer-wise evolution of the spectrum requires further study, however, the matrix consistently tends to uniform attention for large $\ell$.
Scaling discussion.
Our analysis provides practitioners with valuable insights into the scaling of some key quantities in transformer architectures. First, in Figure 8(a), we plot the training loss of a single-layer transformer with key-query attention, for which our theory effectively predicts signal propagation at the initialisation stage. Based on Proposition 1, when the keys $\mathbf{W}^{K}$ and queries $\mathbf{W}^{Q}$ are initialised using Xavier’s initialisation scheme, the attention matrix rapidly converges to degeneracy. Consequently, removing the spectral gap in this case effectively reduces the attention matrix to 00. This explains the plateau in the training curve during the initial steps for the Xavier-initialised model combined with the spectral gap removal. On the other hand, our proposed fix achieves performance similar to that of the original variant. Another scaling that requires refinement is that of a skip connection. Traditionally, the value matrices are sampled from $\mathcal{N}(0,d^{-1})$, the attention matrix is softmax-based, and the signal propagates from one layer $\ell$ to the next as
$\mathbf{X}_{\ell+1}=\mathbf{A}_{\ell+1}(\mathbf{X}_{\ell})\mathbf{X}_{\ell}\mathbf{W}^{V}_{\ell+1}+\mathbf{X}_{\ell}.$
In this scenario, starting from an isometric input $\mathbf{X}_{0}$, the attention mechanism is down-weighted by a factor of $\sqrt{d}$ relative to the skip branch, i.e. the signal coming from the previous layer. Therefore, if the attention mechanism is to be fairly represented in the signal’s propagation through depth, the values should instead be drawn from $\mathcal{N}(0,1)$. If the value weights are sampled from $\mathcal{N}(0,d^{-1})$, the skip connection becomes what we refer to as “upscaled”, meaning each layer essentially passes information from the previous layer without significant transformation. This severely limits the model’s expressivity, reducing its capacity to learn nonlinear mappings. In Figure 8(b), we confirm this observation by comparing the training losses associated with each option. It is not surprising that the “upscaled skip” variants perform worse, as we are attempting to learn a nonlinear function using a virtually linear model.
(a) Xavier initialisation over the keys and queries.
(b) Scaling of the skip connection.
Figure 8: Studying transformers through a spectral lens raises important questions about the validity of some default scaling choices.
@@ 4 Conclusion
We introduced a new mathematical framework for studying the self-attention mechanism at initialisation, leveraging results from random matrix theory and free probability. By analysing the spectral properties of Random Markov matrices, we diagnosed random softmax-based attention with a spectral gap that leads to rank collapse in width—a phenomenon revealed and demonstrated for the first time by our analysis—alongside the previously established rank collapse in depth and exploding gradients.
We proposed a straightforward modification of the attention mechanism, which proved effective in slowing rank collapse when the spectrum contains a single outlier. Additionally, we observed that the spectra of standard key-query attention matrices often feature multiple outliers. Our experiments also pointed to potential issues with some common initialisation schemes for transformers. We hope our work encourages the community to adopt a spectral perspective in investigating more complex transformer architectures and attention models.

### v2

@@ 3.1 Revisiting rank collapse & exploding gradients
Rank collapse
Given an isotropic input $\mathbf{X_{0}}$ with $\boldsymbol{\Sigma}_{0}\coloneqq\mathbf{X}_{0}\mathbf{X}^{\top}_{0}=\mathbf{I}$, we are interested in understanding how the stable rank of the covariance matrix at layer $\ell$,
@@
Proposition 4 (Rank collapse in width for deep network).
@@
with overwhelming probability. Moreover, the convergence happens at a polynomial rate, i.e.
@@
A well-known issue that can disrupt training across various neural network architectures is the vanishing or exploding of gradients; see [Han18]. For attention-only transformers with degenerate attention, [NAB+22] demonstrate that the gradients with respect to $\mathbf{W}^{V}_{\ell}$ vanish. Our model (11) allows for more general random attention while using a different scaling that makes the same quantity explode rather than vanish. Proposition  5 provides a lower bound on the rate at which the gradient grows.
@@
3.2 Removing the gap
As previously seen, we can write an i.i.d. Markov matrix $\mathbf{A}\in\mathbb{R}^{T\times T}$ as
$\mathbf{A}=\frac{1}{T}\mathbf{1}_{T\times T}+\mathbf{A}^{\perp},$ (15)
where $\mathbf{1}_{T\times T}$ is the all-ones matrix and $\mathbf{A}^{\perp}$ has a limiting spectrum resembling that of a Gaussian matrix.
Similar to what we did in the last section to remove the outliers and keep the bulk, we simply replace $\mathbf{A}$ with $\mathbf{A}^{\perp}$ at every layer, i.e.
@@ 3.2 Removing the gap
where $\mathbf{A}^{\perp}_{\ell}\coloneqq\mathbf{A}_{\ell}-T^{-1}\mathbf{1}_{T\times T}$. Note, again, that we directly modify the attention matrices (but not the signal representation) and $\mathbf{X}^{\perp}_{\ell}$ should be understood as the signal at layer $\ell\geq 1$ in a network whose $\mathbf{A}_{\ell}$’s are replaced with $\mathbf{A}^{\perp}_{\ell}$’s as in (16). We set $\mathbf{X}^{\perp}_{0}=\mathbf{X}_{0}$.
The modified attention has no spectral gap (see Lemma 3), thus the stable rank of the covariance matrix $\boldsymbol{\Sigma}^{\perp}_{\ell}\coloneqq\mathbf{X}^{\perp}_{\ell}{\mathbf{X}^{\perp}_{\ell}}^{\top}$ does not collapse anymore, as detailed in Proposition 6 (cf. Proposition 4).
@@
Let $\mathbf{X}^{\perp}_{\ell}=\mathbf{A}^{\perp}_{\ell}\mathbf{X}^{\perp}_{\ell-1}\mathbf{W}^{V}_{\ell}$ be the signal at layer $\ell$ in our modified model (16) and $\mathbf{\Sigma}^{\perp}_{\ell}\in\mathbb{R}^{T\times T}$ be its covariance matrix. Then, almost surely, the rank does not collapse, i.e., there exists a constant $C_{\ell}>0$ such that,
@@
Let $\mathbf{X}^{\perp}_{\ell}=\mathbf{A}^{\perp}_{\ell}\mathbf{X}^{\perp}_{\ell-1}\mathbf{W}^{V}_{\ell}$ be the signal at layer $\ell$ in our modified model (16). Then, in expectation, the squared norm of the gradients grow linearly with $d$, i.e. there exists a constant $C>0$ such that,
@@
4 Experiments and Further Insights
Rank collapse observed in practice.
@@ 4 Experiments and Further Insights
We highlight the practical relevance of our findings by showing rank collapse occurs both in width and depth in widely used transformer encoders like BERT, see Figure 3. In Figure 4(a), all solid lines, corresponding to the vanilla attention layer block, show a stable rank that rapidly converges to $1$ as the context length $T$ grows (at a rate given in Theorem 1). This occurs even if additional components such as LayerNorm, skip connections, or both are added on top of the attention block. Provided that rank collapse in depth is an inherent consequence of repeated matrix multiplications, architectural modifications of the attention layer can only slow the collapse rather than completely prevent it. We demonstrate this by showing that rank collapse in depth persists even when the attention matrix is set to the identity matrix—an extreme case with the highest possible stable rank and no spectral gap.
(a) Rank collapse in width
(b) Rank collapse in depth
Figure 3: Famous transformer encoders suffer from rank collapse at initialisation, both in (a) width and (b) depth. These untrained models are loaded from Hugging Face and sentences from this paper’s abstract are tokenised using pre-trained tokenisers.
Removing the spectral gap.
As an input signal propagates through an initialised transformer, we can address both forms of rank collapse—across width and depth—by eliminating the spectral gap induced by the attention matrix at each layer. Figure 4(a) reinforce our findings, showing that our solution—indicated by in the legend—consistently mitigates rank collapse. Remarkably, the fix is effective even in deep transformers when combined with additional components such as LayerNorm, skip connections, or both. Indeed, the stable rank now scales up with the context length $T$ (at a linear rate predicted by our theory), in contrast to the original behaviour where it would quickly vanish to $1$. Once the rank collapse in width is mitigated, it naturally slows down the one in depth, as seen in Figure 4(b). Another possible way to slow down rank collapse in depth (though not eliminate it) is to set the value matrices as orthogonal matrices.
After passing an isotropic input through the network, we compute the gradient norm as defined in (13). When the attention is an i.i.d. Markov matrix, gradient norms are effectively controlled with depth by either applying LayerNorm or removing the spectral gap, as illustrated in Figure 5. The corresponding lines remain flat, confirming the theoretically expected constant trend with $L$. Without these solutions, the gradients explode exponentially with depth as indicated by the linear trend in the log-log scale, with the derived lower bound of $T^{L-1}$ empirically confirmed.
(a) Spectral gap causes rank collapse in width.
@@
Figure 4: Rank collapse occurs both in width and depth. Our fix effectively prevents the rank from collapsing in width within an attention layer. Although rank collapse in depth occurs regardless of the presence of the spectral gap, our fix consistently slows the collapse—a feat no other module achieves.
Figure 5: In deep transformers initialised with i.i.d. Markov attention, our “remove the gap” fix is effective, as we can precisely discard the single outlier in the spectrum.
Future avenues.
Beyond one attention layer, we conduct an empirical investigation of the spectrum of the actual attention matrix across layers. Surprisingly, as shown in Figure 6, we observe the emergence of additional outliers in the spectrum. We assess the effectiveness of our proposed solution—removing the largest spectral outlier—by comparing it to an alternative approach, which consists in removing all outliers. Since identifying and locating these additional outliers is highly non-trivial, we perform an SVD of the attention matrix at each layer and manually identify and remove the extra gaps based on the largest difference between singular values. As depicted in Figure 7, this computationally expensive approach does not significantly outperform our proposed fix, which is much more efficient. In fact, the two methods achieve comparable results in mitigating rank collapse (first in width, then in depth) and controlling gradient norms. This demonstrates that removing the largest spectral gaps yields the most significant benefit with the lowest cost. Finally, while training dynamics are beyond the scope of this work, our theoretical framework provides a fresh perspective on recent empirical studies that report significant training benefits from various ad hoc fixes. Our analysis offers insights into why these fixes may be effective. These studies include approaches such as “centering” self-attention layers [AGW23], where the authors remove exactly $T^{-1}\mathbf{1}_{T\times T}$ from the attention layer, and “differentiating” the attention mechanism, as seen in the Differential Transformer architecture, where two softmax attention matrices are subtracted from each other at each layer to “cancel the noise”; see [YDX+24]. We can now interpret these approaches as implicit efforts to address the spectral gap.
Figure 6: Spectrum of a $T\times T$ softmax-based attention matrix (with $T=200$) across layers. Some additional outlier eigenvalues emerge from the spectrum with depth in an intricate way. The matrix, however, consistently converges to uniform attention for large $\ell$.
Figure 7: In deep transformers, our “remove the gap” fix is nearly as effective as removing all outliers from the spectrum despite having a significantly lower computational cost.
@@ 5 Conclusion
We introduced a new mathematical framework for studying the self-attention mechanism at initialisation, leveraging results from random matrix theory and free probability. By analysing the spectral properties of i.i.d. Markov matrices, we diagnosed random softmax-based attention with a spectral gap that leads to rank collapse in width—a phenomenon revealed and demonstrated for the first time by our analysis—alongside the previously established rank collapse in depth and exploding gradients.
We proposed a simple modification of the attention mechanism, which is provably effective in slowing rank collapse. Additionally, we empirically discovered that the spectra of standard key-query attention matrices often feature multiple outliers in deeper layers. Context length appears to have been overlooked in previous analyses of rank collapse in favour of depth. Given the continuous growth in the size of transformer models used in practice, we hope this work encourages the community to reconsider the emphasis on depth in theoretical developments, particularly in the analysis of rank collapse.